### v1

@@ Abstract
In the rapidly evolving landscape of autonomous driving, the capability to accurately predict future events and assess their implications is paramount for both safety and efficiency, critically aiding the decision-making process. World models have emerged as a transformative approach, enabling autonomous driving systems to synthesize and interpret vast amounts of sensor data, thereby predicting potential future scenarios and compensating for information gaps. This paper provides an initial review of the current state and prospective advancements of world models in autonomous driving, spanning their theoretical underpinnings, practical applications, and the ongoing research efforts aimed at overcoming existing limitations. Highlighting the significant role of world models in advancing autonomous driving technologies, this survey aspires to serve as a foundational reference for the research community, facilitating swift access to and comprehension of this burgeoning field, and inspiring continued innovation and exploration.
keywords:
[…]
1 Introduction
@@
The quest to develop autonomous driving systems that seamlessly navigate the intricate tapestry of real-world scenarios remains a formidable frontier in contemporary technology. This challenge is not merely technical but also philosophical, probing the essence of cognition and perception that distinguishes human intelligence from artificial constructs. The crux of this challenge lies in imbuing machines with the kind of intuitive reasoning and ‘common sense’ that humans employ effortlessly. Current machine learning systems, despite their prowess, often stumble in pattern recognition tasks that humans resolve with ease, highlighting a significant gap in our quest for truly autonomous systems [1]. On the other hand, human decision-making is deeply rooted in sensory perceptions, constrained by both the memories of these perceptions and direct observations [2, 3]. Beyond mere perception, humans possess the uncanny ability to predict the outcomes of their actions, envision potential futures, and anticipate changes in sensory inputs—abilities that underpin our interaction with the world. The endeavor to replicate such capabilities in machines is not just an engineering challenge but a step towards bridging the cognitive divide between human and machine intelligence.
Addressing this gap, world models have emerged as a critical solution, offering systems the ability to predict and adapt to dynamic environments by emulating human-like perception and decision-making processes. This evolution is essential in the face of real-world scenarios’ complexity and unpredictability, where traditional AI approaches struggle to replicate the depth and variability of human cognitive processes. The necessity for world models is underscored by their potential to bridge the cognitive divide between human and machine intelligence, providing a pathway towards more sophisticated autonomous driving systems.
The journey of world models from conceptual frameworks in control theory during the 1970s to their current prominence in artificial intelligence research reflects a remarkable trajectory of technological evolution and interdisciplinary fusion. The initial formulations in control theory, as laid out by pioneers [4, 5, 6], were foundational, setting the stage for the integration of computational models in dynamic system management. These early efforts were instrumental in demonstrating the potential of applying mathematical models to predict and control complex systems, a principle that would eventually underpin the development of world models.
As the field progressed, the advent of neural networks introduced a paradigm shift, allowing for the modeling of dynamic systems with unparalleled depth and complexity [7, 8, 9]. This transition from static, linear models to dynamic, non-linear representations facilitated a deeper understanding of environmental interactions, laying the groundwork for the sophisticated world models we see today. The integration of recurrent neural networks (RNNs) was particularly transformative, marking a leap towards systems capable of temporal data processing, essential for predicting future states and enabling abstract reasoning [10, 11].
The formal unveiling of world models by Ha and Schmidhuber in 2018 was a defining moment that captured the collective aspiration of the AI research community to endow machines with a level of cognitive processing reminiscent of human consciousness [12, 13]. By harnessing the power of Mixture Density Networks (MDN) and RNNs, this work illuminated the path for unsupervised learning to extract and interpret the spatial and temporal patterns inherent in environmental data. The significance of this breakthrough cannot be overstated, it demonstrated that autonomous systems could achieve a nuanced understanding of their operational environments, predicting future scenarios with an accuracy that was previously unattainable.
@@
Figure 1: Number of Publications related to World Models since 2015. (Data sources: Web of Science Core Collection and Preprint Citation Index.)
@@
In the realm of autonomous driving, the introduction of world models signifies a pivotal shift towards data-driven intelligence, where the capacity to predict and simulate future scenarios becomes a cornerstone for safety and efficiency. The challenge of data scarcity, particularly in specialized tasks such as BEV labeling, underscores the practical necessity of innovative solutions like world models [14, 15, 12]. By generating predictive scenarios from historical data, these models not only circumvent the limitations posed by data collection and labeling but also enhance the training of autonomous systems in simulated environments that can mirror, or even surpass, the complexity of real-world conditions. This approach heralds a new era where autonomous vehicles are equipped with predictive capabilities that reflect a form of intuition, enabling them to navigate and respond to their environment with an unprecedented level of sophistication.
@@
2 Development of World Models
@@
2.1 Architectural Foundations of World Models
@@
1) Perception Module:
This foundational element acts as the system’s sensory input, analogous to human senses. It employs advanced sensors and encoder modules, such as Variational Autoencoder (VAE) [16, 17, 18], Masked Autoencoder (MAE) [19, 20], and Discrete Autoencoder (DAE) [21, 14], to process and compress environmental inputs (images, videos, text, control commands) into a more manageable format. The effectiveness of this module is crucial for the accurate perception of complex, dynamic environments, facilitating a detailed understanding that informs the model’s subsequent predictions and decisions.
2) Memory Module:
Serving a role similar to the human hippocampus, the memory module is pivotal for recording and managing past, present, and predicted world states along with their associated costs or rewards [1]. It supports both short-term and long-term memory functions by replaying recent experiences, a process that enhances learning and adaptation by incorporating past insights into future decisions [22]. This module’s ability to synthesize and retain crucial information is essential for developing a nuanced understanding of environmental dynamics over time.
3) Control/Action Module:
This component is directly responsible for interacting with the environment through actions. It evaluates the current state and predictions provided by the world model to determine the optimal sequence of actions aimed at achieving specific goals, such as minimizing costs or maximizing rewards. The sophistication of this module lies in its ability to integrate sensory data, memory, and predictive insights to make informed, strategic decisions that navigate the complexities of real-world scenarios.
4) World Model Module:
At the heart of the architecture, the world model module performs two primary functions: estimating any missing information about the current world state and predicting future states of the environment. This dual capability allows the system to generate a comprehensive, predictive model of its surroundings, accounting for uncertainties and dynamic changes. By simulating potential future scenarios, this module enables the system to prepare and adjust its strategies proactively, mirroring the predictive and adaptive thought processes found in human cognition.
@@
Figure 2: Diagram of an Agent’s World Model Framework.
In high-dimensional sensory input scenarios, world models utilize latent dynamical models to abstractly represent observed information, enabling compact forward predictions within a latent state space. These latent states, being significantly more space-efficient than direct predictions of high-dimensional data, facilitate the execution of numerous parallel predictions [23], thanks to advancements in deep learning and latent variable models. Take, for instance, the ambiguity of a car’s direction at an intersection, a scenario emblematic of the inherent unpredictability of real-world dynamics. Latent variables serve as a powerful tool in representing these uncertain outcomes, setting the stage for world models to envision a spectrum of future possibilities grounded in the present state. The crux of this endeavor lies in harmonizing the deterministic aspects of prediction with the intrinsic uncertainty of real-world phenomena, a balancing act central to the efficacy of world models.
To navigate this challenge, a variety of strategies have been proposed, from introducing uncertainty via a temperature variable [12] to adopting structured frameworks like the Recurrent State Space Model (RSSM) [23, 24, 25, 26, 27] and the Joint Embedding Predictive Architecture (JEPA) [1, 28, 19, 29]. These methodologies strive to fine-tune the balance between precision and flexibility in prediction. Moreover, leveraging Top-k sampling and transitioning from CNN-based models to transformer architectures [30, 14, 31], such as the Transformer State Space Model (TSSM) or Spatial Temporal Patchwise Transformer (STPT), have shown promise in enhancing model performance by better approximating the complexity and uncertainty of the real world. These solutions strive to align the outputs of world models more closely with the probable developments of the real world. This alignment is crucial because the real world, compared to gaming environments, has a much wider range of influencing factors and a greater degree of randomness in future outcomes. Over-reliance on the highest probability predictions may lead to repetitive cycles in long-term forecasting. Conversely, excessive randomness in predictions can lead to absurd futures that significantly diverge from reality.
In particular, RSSM and JEPA are the most widely used core structures in world model research:
@@
1) Recurrent State Space Model (RSSM) [24] stands as a pivotal model within the Dreamer series of world models, designed to facilitate forward prediction purely within the latent space. This innovative structure enables the model to predict through the latent state space, where both stochastic and deterministic paths within the transition model play critical roles in successful planning.
Figure 3 illustrates a schematic of the latent dynamics model across three time steps. Initially observing two time steps, the model then forecasts the third. Here, stochastic variables (circles) and deterministic variables (squares) interact within the model’s architecture—solid lines illustrate the generative processes, while dashed lines represent inference pathways. The initial deterministic inference approach in Figure 3(a) reveals its limitation in capturing diverse potential futures due to its fixed nature. Conversely, an entirely stochastic approach in Figure 3(b) presents challenges in information retention across time steps, given its inherent unpredictability.
RSSM’s innovation lies in its strategic decomposition of states into stochastic and deterministic components in Figure 3(c), effectively harnessing the predictive stability of deterministic elements and the adaptive potential of stochastic elements. This hybrid structure ensures robust learning and prediction capabilities, accommodating the unpredictability of real-world dynamics while preserving information continuity. By marrying the strengths of RNNs with the flexibility of State Space Models (SSM), RSSM establishes a comprehensive framework for world models, enhancing their ability to predict future states with a balance of precision and adaptability.
We denote the sequence of observations and actions as $\left(\mathbf{x}_{0},\boldsymbol{a}_{1},\mathbf{x}_{1},\boldsymbol{a}_{2},\mathbf{x}_{2},\ldots,\boldsymbol{a}_{T},\mathbf{x}_{T}\right)$. Namely, the agent takes action $\boldsymbol{a}_{t+1}$ after observing $\mathbf{x}_{t}$, and receives the next observation $\mathbf{x}_{t+1}$. We omit the reward for simplicity. RSSM models the observations and state transitions through the following generative process:
$p\left(\mathbf{x}_{0:T}\mid\boldsymbol{a}_{1:T}\right)=\int\prod_{t=0}^{T}p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\boldsymbol{a}_{\leq t}\right)p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t}\right)\mathrm{d}\mathbf{z}_{0:T},$ (1)
@@
$q\left(\mathbf{z}_{0:T}\mid\mathbf{x}_{0:T},\boldsymbol{a}_{1:T}\right)=\prod_{t=0}^{T}q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t},\mathbf{x}_{t}\right)$ (2)
The conditioning on previous states $\mathbf{z}_{<t}$ and actions $\boldsymbol{a}_{\leq t}$ appears multiple times. RSSM uses a shared GRU to compress $\mathbf{z}_{<t}$ and $\boldsymbol{a}_{\leq t}$ into a deterministic encoding $\boldsymbol{h}_{t}$ :
$\boldsymbol{h}_{t}=\operatorname{GRU}\left(\boldsymbol{h}_{t-1},\operatorname{MLP}\left(\operatorname{concat}\left[\mathbf{z}_{t-1},\boldsymbol{a}_{t}\right]\right)\right)$ (3)
@@
$p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t}\right)$ $=\operatorname{MLP}\left(\boldsymbol{h}_{t}\right),$ (4)
$p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\boldsymbol{a}_{\leq t}\right)$ $=\mathcal{N}\left(\hat{\mathbf{x}}_{t},\mathbf{1}\right),\quad\hat{\mathbf{x}}_{t}=\operatorname{Decoder}\left(\operatorname{concat}\left[\boldsymbol{h}_{t},\mathbf{z}_{t}\right]\right),$
$q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t},\mathbf{x}_{t}\right)$ $=\operatorname{MLP}\left(\operatorname{concat}\left[\boldsymbol{h}_{t},\boldsymbol{e}_{t}\right]\right),\quad\boldsymbol{e}_{t}=\operatorname{Encoder}\left(\mathbf{x}_{t}\right).$
@@
$\log p\left(\mathbf{x}_{0:T}\mid\boldsymbol{a}_{1:T}\right)\geq\mathbb{E}_{q}\left[\sum_{t=0}^{T}\log p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\boldsymbol{a}_{\leq t}\right)-\mathcal{L}_{\mathrm{KL}}\left(q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t},\mathbf{x}_{t}\right),p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\boldsymbol{a}_{\leq t}\right)\right)\right]$ (5)
2) Joint-Embedding Predictive Architecture (JEPA) [1] marks a paradigm shift in predictive modeling by focusing on representation space rather than direct, detailed predictions. As shown in Figure 4, by abstracting input ($\mathbf{x}$) and target ($\mathbf{y}$) through dual encoders into representations ($\mathbf{s}_{x}$ and $\mathbf{s}_{y}$), and leveraging a latent variable ($\mathbf{z}$) for prediction, JEPA achieves a significant leap in efficiency and accuracy. This model excels in filtering out noise and irrelevancies, concentrating on the essence of the predictive task. The strategic use of the latent variable ($\mathbf{z}$) to manage uncertainties further refines the model’s focus, enabling it to predict abstract outcomes with heightened precision. By prioritizing relevant features and embracing the inherent uncertainties of predictive tasks, JEPA not only streamlines the prediction process but also ensures outcomes are both relevant and reliable, setting a new standard for predictive modeling in complex environments.
@@
The training of JEPA involve higher-order optimization methods, consider the second-order derivatives to ensure convergence in complex landscapes:
@@
$\log p\left(\mathbf{y}\mid\mathbf{x};\theta,\phi\right)\geq\mathbb{E}_{q\left(\mathbf{z}\mid\mathbf{x};\psi\right)}\left[\log p\left(\mathbf{y}\mid\mathbf{x},z;\theta,\phi\right)\right]-\operatorname{KL}\left[q\left(\mathbf{z}\mid\mathbf{x};\psi\right)\|p\left(\mathbf{z}\mid\mathbf{x};\theta\right)\right]$ (9)

### v2

@@ Abstract
In the rapidly evolving landscape of autonomous driving, the capability to accurately predict future events and assess their implications is paramount for both safety and efficiency, critically aiding the decision-making process. World models have emerged as a transformative approach, enabling autonomous driving systems to synthesize and interpret vast amounts of sensor data, thereby predicting potential future scenarios and compensating for information gaps. This paper provides an initial review of the current state and prospective advancements of world models in autonomous driving, spanning their theoretical underpinnings, practical applications, and ongoing research efforts aimed at overcoming existing limitations. Highlighting the significant role of world models in advancing autonomous driving technologies, this survey aspires to serve as a foundational reference for the research community, facilitating swift access to and comprehension of this burgeoning field, and inspiring continued innovation and exploration.
Index Terms:
World model, Autonomous driving, Foundational model, Model-based reinforcement learning.
I Introduction
@@
The quest to develop autonomous driving systems that seamlessly navigate the intricate tapestry of real-world scenarios remains a formidable frontier in contemporary technology. This challenge is not merely technical but also philosophical, probing the essence of cognition and perception that distinguishes human intelligence from artificial constructs. The crux of this challenge lies in imbuing machines with the kind of intuitive reasoning and ‘common sense’ that humans employ effortlessly. Current machine learning systems, despite their prowess, often stumble in pattern recognition tasks that humans resolve with ease, highlighting a significant gap in our quest for truly autonomous systems [lecun2022path]. On the other hand, human decision-making is deeply rooted in sensory perceptions, constrained by both the memories of these perceptions and direct observations [chang2017code, quiroga2005invariant]. Beyond mere perception, humans possess the uncanny ability to predict the outcomes of their actions, envision potential futures, and anticipate changes in sensory inputs—abilities that underpin our interaction with the world [keller2012sensorimotor]. The endeavor to replicate such capabilities in machines is not just an engineering challenge but a step towards bridging the cognitive divide between human and machine intelligence.
@@
Figure 1: Number of Publications related to World Models since 2015. (Data sources: Web of Science Core Collection and Preprint Citation Index. Key words: ”world model”, ”world models”, ”reinforcement learning”.)
Addressing this gap, world models have emerged as a critical solution, offering systems the ability to predict and adapt to dynamic environments by emulating human-like perception and decision-making processes. This evolution is essential in the face of real-world scenarios’ complexity and unpredictability, where traditional AI approaches struggle to replicate the depth and variability of human cognitive processes. The necessity for world models is underscored by their potential to bridge the cognitive divide between human and machine intelligence, providing a pathway toward more sophisticated autonomous driving systems.
The journey of world models from conceptual frameworks in control theory during the 1970s to their current prominence in artificial intelligence research reflects a remarkable trajectory of technological evolution and interdisciplinary fusion. The initial formulations in control theory, as laid out by pioneers [rault1978model, bryson2018applied, bryson1996optimal], were foundational, setting the stage for the integration of computational models in dynamic system management. The underlying logic of Model Predictive Control (MPC) in control theory aligns closely with that of world models [henaff2019model]. Furthermore, the proposal of the mental model systematically described how humans understand the surrounding world and the relationship between abstract concepts, providing theoretical support for world models to imitate humans’ understanding of the relationship between the world and themselves [forrester1971counterintuitive, johnson2004history, jones2011mental, johnson1983mental]. These early efforts demonstrated the potential of applying mathematical models to predict and control complex systems, providing guidance for modeling human cognitive systems, which ultimately laid the foundation for the development of world models.
As the field progressed, the advent of neural networks introduced a paradigm shift, allowing for the modeling of dynamic systems with unparalleled depth and complexity [kumpati1990identification, levin1993control, draeger1995model]. This transition from static, linear models to dynamic, non-linear representations facilitated a deeper understanding of environmental interactions, laying the groundwork for the sophisticated world models we see today. The integration of recurrent neural networks (RNNs) was particularly transformative, marking a leap towards systems capable of temporal data processing, essential for predicting future states and enabling abstract reasoning [schmidhuber1990line, schmidhuber2015learning].
In the early 1990s, using neural networks to learn models for control and applying predictive models in reinforcement learning for both learning and planning ignited pioneers’ interest in the construction of learnable models, particularly in the contemplation of models’ interactions with the world [narendra1990identification, sutton1991dyna, hunt1992neural, jordan2013forward]. The Dyna architecture elucidated the interaction patterns between agents and the world, as well as the processes of reward and state updates in reinforcement learning [sutton1991dyna]. Research using RNNs to develop internal models for inferring future problems offered a unified framework for the transition from learning to thinking within models [schmidhuber1990making, schmidhuber1990line, chiappa2017recurrent, oh2015action, silver2017predictron, watters2017visual]. Following this, with a long exploration of learning dynamic models and employing these models to train policies, learnable models were back in the center of attention again [deisenroth2011pilco, moerland2023model].
In Judea Pearl’s hierarchy of causation, the base level is SEEING, which involves identifying associations between elements, a stage where most predictive models currently operate; The middle layer, DOING, refers to actions and interventions, exemplified by the exploration of reinforcement learning; The highest level, IMAGINING, focuses on counterfactual reasoning, understanding the causal relationships between things, enabling the transcendence from complete dependence on datasets [pearl2018book]. This capability allows for dealing with situations beyond the training data through imagination. For learnable models, achieving counterfactual reasoning represents an arduous goal [li2024inferring]. While this ability is innate to humans, crossing this threshold poses a significant challenge for AI. Fortunately, the emergence of world models has introduced a turning point in this dilemma.
The formal unveiling of world models by Ha and Schmidhuber in 2018 was a defining moment that captured the collective aspiration of the AI research community to endow machines with a level of cognitive processing reminiscent of human consciousness [ha2018world, ha2018recurrent]. By harnessing the power of Mixture Density Networks (MDN) and RNNs, this work illuminated the path for unsupervised learning to extract and interpret the spatial and temporal patterns inherent in environmental data. The significance of this breakthrough cannot be overstated, it demonstrated that autonomous systems could achieve a nuanced understanding of their operational environments, predicting future scenarios with an accuracy that was previously unattainable.
@@
In the realm of autonomous driving, the introduction of world models signifies a pivotal shift towards data-driven intelligence, where the capacity to predict and simulate future scenarios becomes a cornerstone for safety and efficiency. The challenge of data scarcity, particularly in specialized tasks such as BEV labeling, underscores the practical necessity of innovative solutions like world models [micheli2022transformers, kaiser2019model, ha2018world]. By generating predictive scenarios from historical data, these models not only circumvent the limitations posed by data collection and labeling but also enhance the training of autonomous systems in simulated environments that can mirror, or even surpass, the complexity of real-world conditions. This approach heralds a new era where autonomous vehicles are equipped with predictive capabilities that reflect a form of intuition, enabling them to navigate and respond to their environment with an unprecedented level of sophistication.
@@
II Development of World Models
@@
Figure 2: Diagram of an Agent’s World Model Framework.
II-A Architectural Foundations of World Models
@@
Perception Module
This foundational element acts as the system’s sensory input, analogous to human senses. It employs advanced sensors and encoder modules, such as Variational Autoencoder (VAE) [kingma2013auto, doersch2016tutorial, rezende2014stochastic], Masked Autoencoder (MAE) [bardes2023mc, he2022masked], and Discrete Autoencoder (DAE) [van2017neural, micheli2022transformers], to process and compress environmental inputs (images, videos, text, control commands) into a more manageable format. The effectiveness of this module is crucial for the accurate perception of complex, dynamic environments, facilitating a detailed understanding that informs the model’s subsequent predictions and decisions.
Memory Module
Serving a role similar to the human hippocampus, the memory module is pivotal for recording and managing past, present, and predicted world states along with their associated costs or rewards [lecun2022path]. It supports both short-term and long-term memory functions by replaying recent experiences, a process that enhances learning and adaptation by incorporating past insights into future decisions [foster2017replay]. This module’s ability to synthesize and retain crucial information is essential for developing a nuanced understanding of environmental dynamics over time.
Control/Action Module
This component is directly responsible for interacting with the environment through actions. It evaluates the current state and predictions provided by the world model to determine the optimal sequence of actions aimed at achieving specific goals, such as minimizing costs or maximizing rewards. The sophistication of this module lies in its ability to integrate sensory data, memory, and predictive insights to make informed, strategic decisions that navigate the complexities of real-world scenarios. This module distinguishes the decision-making process from the intricate world model module and trains it independently using a minimal parameter set. This design allows for the application of more unconventional training methods, such as evolution strategies [rechenberg1978evolutionsstrategien], to address challenging reinforcement learning tasks where credit assignment poses significant difficulties.
World Model Module
As the heart of the architecture, the world model module performs two primary functions: estimating any missing information about the current world state and predicting future states of the environment. This dual capability allows the system to generate a comprehensive, predictive model of its surroundings, accounting for uncertainties and dynamic changes. By simulating potential future scenarios, this module enables the system to prepare and adjust its strategies proactively, mirroring the predictive and adaptive thought processes found in human cognition.
@@
In high-dimensional sensory input scenarios, world models utilize latent dynamical models to abstractly represent observed information, enabling compact forward predictions within a latent state space. These latent states, being significantly more space-efficient than direct predictions of high-dimensional data, facilitate the execution of numerous parallel predictions [hafner2019dream], thanks to advancements in deep learning and latent variable models. Take, for instance, the ambiguity of a car’s direction at an intersection, a scenario emblematic of the inherent unpredictability of real-world dynamics. Latent variables serve as a powerful tool in representing these uncertain outcomes, setting the stage for world models to envision a spectrum of future possibilities grounded in the present state. The crux of this endeavor lies in harmonizing the deterministic aspects of prediction with the intrinsic uncertainty of real-world phenomena, a balancing act central to the efficacy of world models [diehl2023uncertainty, diehl2021umbrella].
To navigate this challenge, a variety of strategies have been proposed, from introducing uncertainty via a temperature variable [ha2018world] to adopting structured frameworks like the Recurrent State Space Model (RSSM) [hafner2019dream, hafner2019learning, hafner2020mastering, hafner2023mastering, mendonca2021discovering] and the Joint Embedding Predictive Architecture (JEPA) [lecun2022path, fei2023jepa, bardes2023mc, assran2023self]. These methodologies strive to fine-tune the balance between precision and flexibility in prediction. Moreover, leveraging Top-k sampling and transitioning from CNN-based models to transformer architectures [chen2022transdreamer, micheli2022transformers, zhang2023storm], such as the Transformer State Space Model (TSSM) or Spatial Temporal Patchwise Transformer (STPT), have shown promise in enhancing model performance by better approximating the complexity and uncertainty of the real world. These solutions strive to align the outputs of world models more closely with the probable developments of the real world. This alignment is crucial because the real world, compared to gaming environments, has a much wider range of influencing factors and a greater degree of randomness in future outcomes. Over-reliance on the highest probability predictions may lead to repetitive cycles in long-term forecasting. Conversely, excessive randomness in predictions can lead to absurd futures that significantly diverge from reality.
In particular, the core structures most commonly employed in world model research are RSSM and JEPA:
@@
1) Recurrent State Space Model (RSSM) [hafner2019learning] stands as a pivotal model within the Dreamer series of world models, designed to facilitate forward prediction purely within the latent space. This innovative structure enables the model to predict through the latent state space, where both stochastic and deterministic paths within the transition model play critical roles in successful planning.
Fig. 3 illustrates a schematic of the latent dynamics models across three-time steps. Initially observing two-time steps, these models then forecast the third. Here, stochastic variables (circles) and deterministic variables (squares) interact within the models’ architecture—solid lines illustrate the generative processes, while dashed lines represent inference pathways. The initial deterministic inference approach in Fig. 3a reveals its limitation in capturing diverse potential futures due to its fixed nature. Conversely, an entirely stochastic approach in Fig. 3b presents challenges in information retention across time steps, given its inherent unpredictability.
RSSM’s innovation lies in its strategic decomposition of states into stochastic and deterministic components in Fig. 3c, effectively harnessing the predictive stability of deterministic elements and the adaptive potential of stochastic elements. This hybrid structure ensures robust learning and prediction capabilities, accommodating the unpredictability of real-world dynamics while preserving information continuity. By marrying the strengths of RNNs with the flexibility of State Space Models (SSM) [hamilton1994state], RSSM establishes a comprehensive framework for world models, enhancing their ability to predict future states with a balance of precision and adaptability.
We denote the sequence of observations and actions as $\left(\mathbf{x}_{0},\bm{a}_{1},\mathbf{x}_{1},\bm{a}_{2},\mathbf{x}_{2},\ldots,\bm{a}_{T},\mathbf{x}_{T}\right)$. Namely, the agent takes action $\bm{a}_{t+1}$ after observing $\mathbf{x}_{t}$, and receives the next observation $\mathbf{x}_{t+1}$. We omit the reward for simplicity. RSSM models the observations and state transitions through the following generative process:
$p\left(\mathbf{x}_{0:T}\mid\bm{a}_{1:T}\right)=\int\prod_{t=0}^{T}p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\bm{a}_{\leq t}\right)p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t}\right)\mathrm{d}\mathbf{z}_{0:T}$ (1)
@@
$q\left(\mathbf{z}_{0:T}\mid\mathbf{x}_{0:T},\bm{a}_{1:T}\right)=\prod_{t=0}^{T}q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t},\mathbf{x}_{t}\right)$ (2)
The conditioning on previous states $\mathbf{z}_{<t}$ and actions $\bm{a}_{\leq t}$ appears multiple times. RSSM uses a shared GRU to compress $\mathbf{z}_{<t}$ and $\bm{a}_{\leq t}$ into a deterministic encoding $\bm{h}_{t}$ :
$\bm{h}_{t}=\operatorname{GRU}\left(\bm{h}_{t-1},\operatorname{MLP}\left(\operatorname{concat}\left[\mathbf{z}_{t-1},\bm{a}_{t}\right]\right)\right)$ (3)
@@
$p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t}\right)$ $=\operatorname{MLP}\left(\bm{h}_{t}\right),$ (4)
$p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\bm{a}_{\leq t}\right)$ $=\mathcal{N}\left(\hat{\mathbf{x}}_{t},\mathbf{1}\right),\quad$
$\hat{\mathbf{x}}_{t}$ $=\operatorname{Decoder}\left(\operatorname{concat}\left[\bm{h}_{t},\mathbf{z}_{t}\right]\right),$
$q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t},\mathbf{x}_{t}\right)$ $=\operatorname{MLP}\left(\operatorname{concat}\left[\bm{h}_{t},\bm{e}_{t}\right]\right),$
$\quad\bm{e}_{t}$ $=\operatorname{Encoder}\left(\mathbf{x}_{t}\right).$
@@
$\log p$ $\left(\mathbf{x}_{0:T}\mid\bm{a}_{1:T}\right)\geq\mathbb{E}_{q}\bigg{[}\sum_{t=0}^{T}\log p\left(\mathbf{x}_{t}\mid\mathbf{z}_{\leq t},\bm{a}_{\leq t}\right)$ (5)
$-\mathcal{L}_{\mathrm{KL}}\left(q\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t},\mathbf{x}_{t}\right),p\left(\mathbf{z}_{t}\mid\mathbf{z}_{<t},\bm{a}_{\leq t}\right)\right)\bigg{]}$
(a) Joint-Embedding Architecture
(b) Generative Architecture
(c) Joint-Embedding Predictive Architecture
2) Joint-Embedding Predictive Architecture (JEPA) [lecun2022path] marks a paradigm shift in predictive modeling by focusing on representation space rather than direct, detailed predictions. As shown in Fig. 4, by abstracting input ($\mathbf{x}$) and target ($\mathbf{y}$) through dual encoders into representations ($\mathbf{s}_{x}$ and $\mathbf{s}_{y}$), and leveraging a latent variable ($\mathbf{z}$) for prediction, JEPA achieves a significant leap in efficiency and accuracy. This model excels in filtering out noise and irrelevancies, concentrating on the essence of the predictive task. The strategic use of the latent variable ($\mathbf{z}$) to manage uncertainties further refines the model’s focus, enabling it to predict abstract outcomes with heightened precision. By prioritizing relevant features and embracing the inherent uncertainties of predictive tasks, JEPA not only streamlines the prediction process but also ensures outcomes are both relevant and reliable, setting a new standard for predictive modeling in complex environments.
@@
The training of JEPA involves higher-order optimization methods, considering the second-order derivatives to ensure convergence in complex landscapes:
@@
$\log p\left(\mathbf{y}\mid\mathbf{x};\theta,\phi\right)\geq\mathbb{E}_{q\left(\mathbf{z}\mid\mathbf{x};\psi\right)}\left[\log p\left(\mathbf{y}\mid\mathbf{x},z;\theta,\phi\right)\right]$ (9)
$-\operatorname{KL}\left[q\left(\mathbf{z}\mid\mathbf{x};\psi\right)\|p\left(\mathbf{z}\mid\mathbf{x};\theta\right)\right]$